# Towards Pervasive Distributed Agentic Generative AI - A State of The Art

### v1

@@ 3.2.2. Fog Computing
  • Device and Network Management: Due to its decentralised and heterogeneous nature, managing randomly distributed network resources and a high-risk device breakdowns complicate connectivity and application deployment.
@@
  • Computational Challenges: The hierarchical structure of fog systems and their interaction with the cloud make optimal task allocation across IoT devices, fog nodes, and the cloud challenging. Ensuring computational correctness in such distributed environments is complex, further compounded by the need to choose suitable protocols for heterogeneous sensors and devices.
  • Security Challenges: Heterogeneous devices, in less secure environments, can be susceptible to various attacks, including man-in-the-middle attacks.
3.2.3. Edge Computing
Sometimes referred to as an IoT network, performs computations closer to the edge of the network, where the data is generated. As defined by the National Institute of Standards and Technology (NIST), edge computing ’is the network layer including the end-devices and their users, to provide, for example, local computing capability on a sensor, metering or some other devices that are network-accessible’ (Iorga et al., 2017). Decentralizing intelligence, processing power, and communication resources to the network edge, this paradigm leverages a heterogeneous range of devices (including PCs, smartphones, smart devices, and automation controllers) situated in close proximity to the data-generating entity, thereby enhancing responsiveness and efficiency (Yalli et al., 2024). Furthermore, processing sensitive data locally on edge devices enhances user privacy and data ownership by minimizing the need for data transfer to the cloud. This localised processing is particularly advantageous for applications dealing with personal or proprietary information, such as virtual assistants and autonomous vehicles. Mobile Edge Computing (MEC) deploys services and computational capabilities at the edge of cellular networks, in close proximity to subscribers. This strategic placement enables a service environment characterised by ultra-low latency, high bandwidth, and direct access to real-time network information (Li et al., 2016). For instance in (Sun and Ansari, 2016), the researchers handle data streams at the mobile edge to overcome the scalability problem of traditional IoT architectures, reducing traffic load in the core network and end-to-end delay for IoT services. The implementation and management of edge computing environments are associated with several challenges, notably resource and energy limitations. Individual edge devices inherently possess constrained computational power, memory capacity, and energy resources compared to centralised cloud servers, necessitating careful consideration on efficiency.
3.3. Pervasive AI Computing
@@
The idea is to use the capabilities of AI systems inside a pervasive computing environment. As defined in (Baccour et al., 2022) pervasive AI is the ’intelligent and efficient distribution of AI tasks and models over/among any types of devices with heterogeneous capabilities in order to execute sophisticated global missions’. This paradigm marks a departure from traditional, cloud centralised AI approaches, leveraging the distributed computational resources inherent in pervasive environments, including IoT devices and edge servers. Within this domain, AI-enabled sensors play a crucial role, categorised as: AIoT Sensors, facilitating cloud-based AI decision-making based on physical-world data; Edge AI Sensors, enabling localised AI inference at the device level; and TinyML Sensors, designed for efficient execution of specific tasks with minimal data (Sloane et al., 2025). The convergence of pervasive computing and artificial intelligence has thus established a novel research area, significantly enabled by advancements in Deep Learning.
@@ 3.3.1. Deep Learning
  • Distributed inference: Splits DL models across devices to reduce local load and cloud latency, e.g. EdgeShard (Zhang et al., 2024f).
  • Federated learning (FL): Trains models across devices while keeping data local, supporting privacy and scalability (Qu et al., 2025).
  • Optimization techniques: Include caching, compression, and dynamic inference to improve DNN performance on constrained devices (Siam et al., 2025).
Within this context, Large Language Models (LLMs), are increasingly demonstrating their potential. When integrated into agent-based systems, they show strong potential for enhancing pervasive AI applications. The next section explores how LLM-based agents are applied in pervasive computing and the challenges of deploying them on limited-resource platforms.
@@ 4. Agents in Pervasive Computing
The integration of intelligent agents within pervasive computing environments represents a significant advancement in human-computer interaction, leading to more intuitive and proactive experiences. Pervasive computing, characterised by the integration of embedding of computational capabilities into everyday objects and environments, provides an ideal environment for the deployment of agents capable of perceiving, reasoning, and acting autonomously. As discussed in Section 2, the advent of Large Language Models (LLMs) has substantially augmented the potential of these agents by equipping them with advanced natural language understanding and generation capabilities, enabling their application in a wide range of use cases. This enhancement facilitates and increase the performance of this pervasive devices with more intuitive and natural interactions within smart environments, such as automatic voice-controlled smart homes (Rivkin et al., 2024), automatic smartphone interactions (Wen et al., 2024a), or automatic traffic management in smart cities (Chen et al., 2024b). However, the deployment of LLM-powered agents in pervasive computing scenarios presents various challenges. As discussed in Section 3, one of the primary concern arises from the resource constraints inherent in many pervasive devices, including limitations in processing power and energy availability. This contrasts with the considerable computational, energy, and memory storage demands of LLMs. Moreover, the direct interaction of these agents with user data on pervasive devices underscores the critical importance of user data privacy. Consequently, ensuring both the sustainable operation of these agents and the robust safeguarding of user data are essential for their responsible development and deployment. Addressing these challenges necessitates research into several key areas. This includes determining optimal deployment and alignment strategies for Large Language Models in both local and distributed pervasive environments. Also, research must focus on redesigning the core components of agents specifically memory, planning, reasoning, and acting modules to operate efficiently within the resource limitations inherent in pervasive computing. Furthermore, rigorous evaluation of agent performance within these specific pervasive environments is crucial to identify their strengths and weaknesses.
@@ 4.1. LLMs Deployment Strategies
Addressing the integration of Large Language Models (LLMs) into resource-constrained pervasive computing environments involves in two primary deployment strategies: local and distributed. The local deployment strategy aims to execute the LLM directly on the device where the user data is stored, minimizing latency and enhancing data privacy. However, this approach is constrained by the limited computational and resource capabilities of the local device. The distributed deployment strategy aims to augment computational resources through cloud computing servers, which offer enhanced processing power but introduce greater latency and raise privacy issues. Alternatively, this strategy involves partitioning the Large Language Model across multiple fog or edge devices, a method that strategically locates processing closer to the user’s device to decrease latency, but increase system complexity.
@@ 4.1.1. Local LLM Deployment
Local deployment of LLMs aims to execute these powerful models directly on resource-constrained edge devices, offering significant advantages in terms of user privacy, cost-effectiveness by eliminating reliance on cloud infrastructure, and reduced latency due to on-device processing. However, the inherent limitations in computational power, energy availability, and storage capacity of these pervasive devices present substantial challenges to implement this vision. A key strategy to overcome these limitations involves optimizing LLM resource consumption through techniques such as small language models (SLMs) deployment, model quantization (to reduce the bit-precision of model weights), and pruning techniques (to reduce model parameters). For instance, researchers in (Wen et al., 2024a) successfully implemented an agentic architecture using a Vicuna-7B, on a OnePlus ACE 2 Pro smartphone equipped with a Snapdragon 8 Gen2 CPU and Adreno™ 740 GPU. Building upon this, Autodroid v2 (Wen et al., 2024b) demonstrated impressive performance in success rate and inference latency on the same resource-constrained device by deploying an 8-bit quantised Llama3.1-8B model. Further highlighting the potential of smaller models, Microsoft’s Magma (Yang et al., 2025a), an 8B parameter model, exhibited strong verbal and spatial-temporal reasoning in UI navigation and robotic manipulation tasks. Additionally, the MaViLa framework (Fan et al., 2025) showcased the effectiveness of smaller models by fine-tuning a Vicuna-13B model for smart manufacturing, achieving high performance in domain-specific tasks like additive manufacturing monitoring, anomaly detection, and autonomous process optimization. Despite the relative smallness of 7B or 13B parameter models compared to their larger 30B+ counterparts, their deployment on resource-constrained devices like typical smartphones (with 6-12 GB of RAM) remains a significant obstacle. As previously discussed, the feasibility of local LLM execution is heavily influenced by the model’s parameter count and the memory footprint of the Key-Value (KV) cache. A Llama-2 7B model, for example, can require up to 28 GB for full precision inference, 7 GB with 8-bit quantization, and 3.5 GB with even 4-bit quantised versions, in addition to over 2 GB potentially needed for its 4k token context window’s KV cache. While deployment might be viable on more powerful edge devices like NVIDIA Jetson Orin series  [^2] Jetson Orin Technical Specification for developer and industrial applications, running such models efficiently alongside their KV cache on devices with less computational power, necessitates innovative memory management strategies. Furthermore, energy consumption is a critical concern, with research indicating an approximate cost of 0.1 J/token per billion parameters (Liu et al., 2024b). This suggests that a 7B parameter LLM could consume 0.7 J/token, potentially limiting continuous conversational usage. For instance a fully charged iPhone, with approximately 50 kJ of energy, can sustain this model in conversation for less than 2 hours at a rate of 10 tokens/s, with every 64 tokens draining 0.2% of the battery (Liu et al., 2024b). This highlights the urgent need for techniques that significantly enhance the energy efficiency of these models. To address these challenges, the literature presents various promising models and methodologies. Nexa AI’s OmniVLM (Chen et al., 2024a), a compact model with under one billion parameters (968M), directly tackles the memory footprint and computational demand issues, making it more easy to deploy. OmniVLM also introduces a novel token compression mechanism for visual inputs, achieving a substantial reduction in visual token sequence length, thereby lowering computational overhead while preserving visual-semantic information. Octopus v2 (Chen and Li, 2024a) uses another approach, enabling a 2 billion parameter on-device language model to outperform GPT-4 in function calling accuracy and latency while drastically reducing context length. This significantly enhances the feasibility of deploying AI agents directly on edge devices without cloud reliance. Building on this success, Octopus v3 (Chen and Li, 2024b) introduces a sub-billion parameter multimodal model capable of efficiently processing both visual and textual inputs using functional tokens and CLIP-based image encoding. Beyond model optimization, innovative deployment and runtime techniques are being explored. EdgeLLM (Yu et al., 2024) proposes a layerwise unified compression (LUC) method for dynamic pruning and quantization of LLM layers, coupled with adaptive layer tuning and voting. This approach achieves significant reductions in computation and memory demands, enabling fine-tuning and updating of LLMs even on smartphones with substantial speedups and reduced memory usage. Another strategy, explored in (Yin et al., 2024), involves hosting a single, stateful LLM as a system service within the mobile operating system, accessible to applications via system APIs. This design minimizes memory duplication and supports persistent context management through chunk-wise KV cache compression and tolerance-aware memory management. Context switching is accelerated via a swapping-recompute pipeline that overlaps I/O and computation. The system employs fine-grained memory eviction strategies, such as LCTRU (Least Compression-Tolerable and Recently-Used queue), adapted to the compression sensitivity of different parts of the model. Finally, EdgeMoE (Yi et al., 2025) faces the challenge of deploying extremely large and sparse Mixture-of-Experts (MoE) LLMs on mobile hardware. By treating device memory as a smart cache and selectively preloading only the most likely needed experts, EdgeMoE can execute models with over 10 billion parameters on small edge devices with minimal overhead. Its key innovations include expert-wise bitwidth adaptation and predictive expert caching, enabling real-time inference without exhausting device resources.
@@ 4.1.2. Distributed LLM Deployment
Cloud computing offers the capability to host large open-source or proprietary LLMs on powerful servers, often accessed via Web Applications or APIs provided by major companies such as OpenAI (offering models like GPT-4, GPT-4o, o1, and o3), Anthropic (providing models like Claude 2.1, Claude 3.5 Sonnet, Claude 3.5 Haiku, and Claude 3.7), or Mistral(providing models like Mistral Large and Mistral Small). In this architecture, edge devices transmit user prompts to these remote servers hosting the LLMs. The server receives the prompt, processes it with the designated model, generates a response, and then sends the answer back to the originating edge device. Cloud computing proves advantageous when high-performance LLMs are required. However, as the servers and models are typically managed by external entities, careful consideration must be given to user privacy, the potential for connection instability, and response latency. Examples of cloud-based LLM applications include (Rivkin et al., 2024), where models like GPT-4 and Claude 2.1 are employed for reasoning and managing smart home devices; MobileGPT (Lee et al., 2024) on smartphones, which uses GPT-4 to manage mobile applications; and Intrusion Detection System Agent (Li et al., 2024b), where GPT-4o reasons over network traffic data, generating and executing actions such as data preprocessing, classification, and knowledge retrieval for intrusion detection with detailed explanations. In scenarios where a single LLM may exhibit limitations across diverse domains, a multi-LLM approach can be employed. This involves using multiple LLMs from different cloud providers, managed by a router. Upon receiving a user request, the router intelligently directs the query to the model best specialised for that specific domain, optimizing the response quality. For instance, Nexa AI’s Octopus v4 (Chen and Li, 2024c) is an LLM agent designed to run locally on the user’s device. It analyzes incoming requests and determines the most appropriate model to handle them, preparing the prompt to maximise the chosen model’s performance. This agent can route calls to both cloud-hosted and edge-hosted LLMs. Additionally, Division-of-Thoughts (DoT) (Shao et al., 2025) is a framework that combines SLMs with powerful cloud LLMs to efficiently handle complex tasks. It first decomposes a user’s query into simpler sub-tasks using a Task Decomposer, exploiting the reasoning abilities of language models. A Task Scheduler then analyzes dependencies among sub-tasks to decide which ones can be executed locally and which need cloud support. A lightweight, plug-and-play Adapter helps the SLM decide task allocation without changing its core parameters. This collaboration reduces costs, boosts speed, and preserves reasoning quality.
@@
Edge and fog computing deploys models closer to users to minimise communication latency, while simultaneously maintaining significant computational power by distributing workloads across multiple devices. This approach enhances user privacy and system modularity, as models operate within user-managed nodes, ultimately improving overall system responsiveness. Although lacking the computational power of cloud environments, these distributed architectures effectively harness the combined resources of numerous edge or fog nodes to distribute computation. This collaborative use of resources enables the deployment of larger, more capable models and the achievement of faster inference times compared to the constraints of a single edge device, as previously outlined. However, a significant challenge lies in the complexity of managing and maintaining the distributed nodes and orchestrating the computation across them, requiring considerable effort from the user. The literature presents several methodologies to address this distributed deployment paradigm. Some strategies focus on dynamic resource allocation and intelligent task distribution such as SpeziLLM (Zagar et al., 2025) and Ai Flow (Shao and Li, 2025). SpeziLLM is an open-source framework that dynamically distributes LLM inference across decentralised fog and edge layers, in healthcare applications. By abstracting orchestration tasks like node selection, model placement, and task splitting, SpeziLLM simplifies the integration of LLMs into mobile and healthcare environments. It prioritizes the execution of sensitive data tasks on trusted local or fog nodes while offloading less critical or computationally intensive tasks to the cloud when necessary, balancing privacy, cost, and user experience through seamless model migration, fault tolerance, and flexible scaling. AI Flow redefines communication by focusing on ”intelligence flow” rather than raw data transfer, adapting to dynamic network conditions to optimise inference across devices, edge nodes, and cloud servers. Instead of transmitting raw data, AI Flow sends only critical extracted features, significantly reducing communication overhead. It adaptively assigns portions of the inference task based on available computational resources, network bandwidth, and real-time conditions, ensuring low-latency responses and efficient model execution even in fluctuating environments. Other strategies work on collaborative inference among edge devices such as Distributed Mixture-of-Agents (MoA) (Mitra et al., 2024). This architecture enables multiple edge devices, each hosting a localised LLM, to collaborate through decentralised gossip protocols, achieving high-quality responses without a centralised server. Each device can independently process prompts and share intermediate results with neighboring devices using decentralized gossip algorithms. Devices act as ”proposers” generating answers and ”aggregators” refining or selecting the best response. This distributed setup enhances robustness, reduces latency, and improves answer quality compared to relying on a single device, while also ensuring queue stability despite resource limitations and varying workloads. Another approach is using the edge devices to model partitioning and distribution, cooperating as a single high performance edge device. EdgeShard (Zhang et al., 2024f) is a method that facilitates efficient LLM inference by partitioning large models into smaller ”shards” and distributing them across multiple edge devices. It intelligently selects devices and allocates model parts based on their computing power, memory, and network conditions, leveraging collaborative edge computing to reduce latency, bandwidth usage, and privacy risks. By employing dynamic programming algorithms for optimized device selection and task scheduling, EdgeShard enables large models like Llama2-70B to run efficiently even in heterogeneous, resource-limited environments, significantly improving inference speed and throughput without compromising model accuracy. In (Zhao et al., 2024b) the edges are used for a cooperative inference with terminal devices. In this framework efficient LLM inference is enabled by promoting collaboration between the user’s device (terminal) and a nearby edge server. The terminal device rapidly generates speculative tokens using a lightweight model, while the edge server concurrently verifies and corrects them using a larger, more accurate LLM. This serial-parallel approach significantly reduces token generation delay and energy consumption by balancing the computational load between local and edge resources without heavy reliance on the cloud. An optimization algorithm manages model approximation and token generation to minimise delay and maintain high accuracy. Finally in distributed scenarios, guaranteeing service despite intermittent connectivity is crucial. In (Zimmer et al., 2024) the authors address this problem with a novel ”Mixture of Attentions” architecture for speculative decoding. This method significantly enhances a local small model’s autonomous prediction by integrating Layer Self-Attention and Cross-Attention, effectively using LLM activations when available, yet maintaining accuracy when disconnected. This approach boosts robustness, enabling continuous, accurate inference without persistent network access.

### v2

@@ 3.1.2. Edge Layer
  • Device and Network Management: Due to its decentralised and heterogeneous nature, managing randomly distributed network resources and high-risk device breakdowns complicates connectivity and application deployment.
@@
  • Computational Challenges: The hierarchical structure of edge systems makes optimal task allocation across IoT devices, intermediate nodes, and the cloud challenging. Ensuring computational correctness in such distributed environments is complex, further compounded by the need to choose suitable protocols for heterogeneous sensors and devices.
  • Security Challenges: Heterogeneous devices in less secure physical environments can be susceptible to various attacks, including man-in-the-middle attacks. Furthermore, while processing sensitive data locally enhances privacy by minimizing cloud transfer, it requires robust local security measures to maintain user trust.
4. Pervasive AI Computing
@@
The integration of Artificial Intelligence into the pervasive computing domain is driven by its strong capability to resolve complex tasks that traditional computational methods cannot handle efficiently. From classifying intricate patterns in images and text to generating sophisticated sequences of actions, AI transforms various heterogeneous environmental inputs into actionable intelligence. This integration is significantly enabled by the rapid advancements in computational hardware, which allow the deployment of AI models even on resource constrained devices. This convergence of advanced problem-solving capabilities and hardware evolution has established the field of Pervasive AI (baccour2022pervasive) defined as the ’intelligent and efficient distribution of AI tasks and models over/among any types of devices with heterogeneous capabilities in order to execute sophisticated global missions’. This paradigm marks a departure from traditional, cloud centralised AI approaches, leveraging the distributed computational resources inherent in pervasive environments, including IoT devices and edge servers. Within this domain, AI-enabled sensors play a crucial role, categorised as: AIoT Sensors, facilitating cloud-based AI decision-making based on physical-world data; Edge AI Sensors, enabling localised AI inference at the device level; and TinyML Sensors, designed for efficient execution of specific tasks with minimal data (sloane2025materiality). In this field, Deep Learning plays a key role in the diffusion and application of AI models within pervasive computing.
@@ 4.0.1. Deep Learning
  • Distributed inference: Splits DL models across devices to reduce local load and cloud latency, e.g. EdgeShard (zhang2024edgeshard).
  • Federated learning (FL): Trains models across devices while keeping data local, supporting privacy and scalability (qu2025mobile).
  • Optimization techniques: Include caching, compression, and dynamic inference to improve DNN performance on constrained devices (siam2025artificial).
While traditional DL models have empowered pervasive devices to effectively perceive complex sensory data (e.g., computer vision, text processing), they operate as rigid inference engines. These models typically depend on strictly structured inputs and are confined to a pre-determined set of outputs, effectively mapping specific signals to fixed labels or commands. The novelty of Agentic AI lies in the integration of Large Language Models (LLMs), which fundamentally disrupts this paradigm. Unlike traditional models, LLMs can interpret complex, unstructured natural language queries without requiring a rigid input format. They possess the ability to generalize user intent, infer logical plans, and generate dynamic action sequences to resolve novel tasks. This shift advances pervasive systems from merely classifying environmental states to actively reasoning about them, thereby enabling autonomous problem-solving in dynamic environments.
The next section explores how these LLM-based agents are specifically applied in pervasive computing and addresses the significant challenges of deploying such powerful reasoning engines on limited-resource platforms.
@@ 5. Agents in Pervasive Computing
The integration of intelligent agents within pervasive computing environments represents a significant advancement in human-computer interaction, leading to more intuitive and proactive experiences. Pervasive computing, characterised by the integration of embedding of computational capabilities into everyday objects and environments, provides an ideal environment for the deployment of agents capable of perceiving, reasoning, and acting autonomously. As discussed in Section 2, the advent of Large Language Models (LLMs) has substantially augmented the potential of these agents by equipping them with advanced natural language understanding and generation capabilities, enabling their application in a wide range of use cases. This enhancement facilitates and increase the performance of this pervasive devices with more intuitive and natural interactions within smart environments, such as automatic voice-controlled smart homes (rivkin2024aiot), automatic smartphone interactions (wen2024autodroidv1), or automatic traffic management in smart cities (chen2024llm). However, the deployment of LLM-powered agents in pervasive computing scenarios presents various challenges. As discussed in Section 3, one of the primary concern arises from the resource constraints inherent in many pervasive devices, including limitations in processing power and energy availability. This contrasts with the considerable computational, energy, and memory storage demands of LLMs. Moreover, the direct interaction of these agents with user data on pervasive devices underscores the critical importance of user data privacy. Consequently, ensuring both the sustainable operation of these agents and the robust safeguarding of user data are essential for their responsible development and deployment. Addressing these challenges necessitates research into several key areas. This includes determining optimal deployment and alignment strategies for LLMs in both local and distributed pervasive environments. Also, research must focus on redesigning the core components of agents specifically memory, planning, reasoning, and acting modules to operate efficiently within the resource limitations inherent in pervasive computing. Furthermore, rigorous evaluation of agent performance within these specific pervasive environments is crucial to identify their strengths and weaknesses.
To address these challenges, this section provides a comprehensive analysis of the adaptation of agents for pervasive environments. First, we examine the deployment strategies of the LLMs, evaluating the trade-offs between local execution and distributed (edge/cloud) architectures. Next, we discuss the alignment strategies, focusing on techniques like quantization and fine-tuning that tailor models to resource-constrained domains. We then explore the agent design strategies, detailing how core the core modules (memory, planning, and action) are adapted for pervasive constraints. Finally, we review real-world evaluation metrics and applications across various pervasive environments such as smart homes, mobile devices, and healthcare.
@@ 5.1. LLMs Deployment Strategies
Addressing the integration of LLMs into resource-constrained pervasive computing environments involves in two primary deployment strategies: local and distributed. The local deployment strategy aims to execute the LLM directly on the device where the user data is stored, minimizing latency and enhancing data privacy. However, this approach is constrained by the limited computational and resource capabilities of the local device. The distributed deployment strategy aims to augment computational resources through cloud computing servers, which offer enhanced processing power but introduce greater latency and raise privacy issues. Alternatively, this strategy involves partitioning the Large Language Model across multiple fog or edge devices, a method that strategically locates processing closer to the user’s device to decrease latency, but increases system complexity.
@@ 5.1.1. Local LLM Deployment
Local deployment of LLMs aims to execute these powerful models directly on resource-constrained edge devices, offering significant advantages in terms of user privacy, cost-effectiveness by eliminating reliance on cloud infrastructure, and reduced latency due to on-device processing. However, the inherent limitations in computational power, energy availability, and storage capacity of these pervasive devices present substantial challenges to implement this vision. A key strategy to overcome these limitations involves optimizing LLM resource consumption through techniques such as small language models (SLMs) deployment, model quantization (to reduce the bit-precision of model weights), and pruning techniques (to reduce model parameters). For instance, researchers in (wen2024autodroidv1) successfully implemented an agentic architecture using a Vicuna-7B, on a OnePlus ACE 2 Pro smartphone equipped with a Snapdragon 8 Gen2 CPU and Adreno™ 740 GPU. Building upon this, Autodroid v2 (wen2024autodroid) demonstrated impressive performance in success rate and inference latency on the same resource-constrained device by deploying an 8-bit quantised Llama3.1-8B model. Further highlighting the potential of smaller models, Microsoft’s Magma (yang2025magma), an 8B (where ’B’ denotes Billion) parameter model, exhibited strong verbal and spatial-temporal reasoning in UI navigation and robotic manipulation tasks. Additionally, the MaViLa framework (fan2025mavila) showcased the effectiveness of smaller models by fine-tuning a Vicuna-13B model for smart manufacturing, achieving high performance in domain-specific tasks like additive manufacturing monitoring, anomaly detection, and autonomous process optimization. Despite the relative smallness of 7B or 13B parameter models compared to their larger 30B+ counterparts, their deployment on resource-constrained devices like typical smartphones (with 6-12 GB of RAM) remains a significant obstacle. As previously discussed, the feasibility of local LLM execution is heavily influenced by the model’s parameter count and the memory footprint of the Key-Value (KV) cache. A Llama-2 7B model, for example, can require up to 28 GB for full precision inference, 7 GB with 8-bit quantization, and 3.5 GB with even 4-bit quantised versions, in addition to over 2 GB potentially needed for its 4k token context window’s KV cache. While deployment might be viable on more powerful edge devices like NVIDIA Jetson Orin series  [^2] Jetson Orin Technical Specification for developer and industrial applications, running such models efficiently alongside their KV cache on devices with less computational power, necessitates innovative memory management strategies. Furthermore, energy consumption is a critical concern, with research indicating an approximate cost of 0.1 J/token per billion parameters (liu2024mobilellm). This suggests that a 7B parameter LLM could consume 0.7 J/token, potentially limiting continuous conversational usage. For instance a fully charged iPhone, with approximately 50 kJ of energy, can sustain this model in conversation for less than 2 hours at a rate of 10 tokens/s, with every 64 tokens draining 0.2% of the battery (liu2024mobilellm). This highlights the urgent need for techniques that significantly enhance the energy efficiency of these models. To address these challenges, the literature presents various promising models and methodologies. Nexa AI’s OmniVLM (chen2024omnivlm), a compact model with under one billion parameters (968M), directly tackles the memory footprint and computational demand issues, making it more easy to deploy. OmniVLM also introduces a novel token compression mechanism for visual inputs, achieving a substantial reduction in visual token sequence length, thereby lowering computational overhead while preserving visual-semantic information. Octopus v2 (chen2024octopusv2) uses another approach, enabling a 2 billion parameter on-device language model to outperform GPT-4 in function calling accuracy and latency while drastically reducing context length. This significantly enhances the feasibility of deploying AI agents directly on edge devices without cloud reliance. Building on this success, Octopus v3 (chen2024octopusv3) introduces a sub-billion parameter multimodal model capable of efficiently processing both visual and textual inputs using functional tokens and CLIP-based image encoding. Beyond model optimization, innovative deployment and runtime techniques are being explored. EdgeLLM (yu2024edge) proposes a layerwise unified compression (LUC) method for dynamic pruning and quantization of LLM layers, coupled with adaptive layer tuning and voting. This approach achieves significant reductions in computation and memory demands, enabling fine-tuning and updating of LLMs even on smartphones with substantial speedups and reduced memory usage. Another strategy, explored in (yin2024llm), involves hosting a single, stateful LLM as a system service within the mobile operating system, accessible to applications via system APIs. This design minimizes memory duplication and supports persistent context management through chunk-wise KV cache compression and tolerance-aware memory management. Context switching is accelerated via a swapping-recompute pipeline that overlaps I/O and computation. The system employs fine-grained memory eviction strategies, such as LCTRU (Least Compression-Tolerable and Recently-Used queue), adapted to the compression sensitivity of different parts of the model. Finally, EdgeMoE (yi2025edgemoe) faces the challenge of deploying extremely large and sparse Mixture-of-Experts (MoE) LLMs on mobile hardware. By treating device memory as a smart cache and selectively preloading only the most likely needed experts, EdgeMoE can execute models with over 10 billion parameters on small edge devices with minimal overhead. Its key innovations include expert-wise bitwidth adaptation and predictive expert caching, enabling real-time inference without exhausting device resources.
@@ 5.1.2. Distributed LLM Deployment
Cloud computing offers the capability to host large open-source or proprietary LLMs on powerful servers, often accessed via Web Applications or APIs provided by major companies such as OpenAI (offering models like GPT-4, GPT-4o, o1, and o3), Anthropic (providing models like Claude 2.1, Claude 3.5 Sonnet, Claude 3.5 Haiku, and Claude 3.7), or Mistral (providing models like Mistral Large and Mistral Small). In this architecture, edge devices transmit user prompts to these remote servers hosting the LLMs. The server receives the prompt, processes it with the designated model, generates a response, and then sends the answer back to the originating edge device. Cloud computing proves advantageous when high-performance LLMs are required. However, as the servers and models are typically managed by external entities, careful consideration must be given to user privacy, the potential for connection instability, and response latency. Examples of cloud-based LLM applications include (rivkin2024aiot), where models like GPT-4 and Claude 2.1 are employed for reasoning and managing smart home devices; MobileGPT (lee2024mobilegpt) on smartphones, which uses GPT-4 to manage mobile applications; and Intrusion Detection System Agent (li2024ids), where GPT-4o reasons over network traffic data, generating and executing actions such as data preprocessing, classification, and knowledge retrieval for intrusion detection with detailed explanations. In scenarios where a single LLM may exhibit limitations across diverse domains, a multi-LLM approach can be employed. This involves using multiple LLMs from different cloud providers, dispatched by a router. Upon receiving a user request, the router intelligently directs the query to the model best specialised for that specific domain, optimizing the response quality. For instance, Nexa AI’s Octopus v4 (chen2024octopusv4) is an LLM agent designed to run locally on the user’s device. It analyzes incoming requests and determines the most appropriate model to handle them, preparing the prompt to maximise the chosen model’s performance. This agent can route calls to both cloud-hosted and edge-hosted LLMs. Additionally, Division-of-Thoughts (DoT) (shao2025division) is a framework that combines SLMs with powerful cloud LLMs to efficiently handle complex tasks. It first decomposes a user’s query into simpler sub-tasks using a Task Decomposer, exploiting the reasoning abilities of language models. A Task Scheduler then analyzes dependencies among sub-tasks to decide which ones can be executed locally and which need cloud support. A lightweight, plug-and-play Adapter helps the SLM decide task allocation without changing its core parameters. This collaboration reduces costs, boosts speed, and preserves reasoning quality.
@@
Edge and fog computing deploys models closer to users to minimise communication latency, while simultaneously maintaining significant computational power by distributing workloads across multiple devices. This approach enhances user privacy and system modularity, as models operate within user-managed nodes, ultimately improving overall system responsiveness. Although lacking the computational power of cloud environments, these distributed architectures effectively harness the combined resources of numerous edge or fog nodes to distribute computation. This collaborative use of resources enables the deployment of larger, more capable models and the achievement of faster inference times compared to the constraints of a single edge device, as previously outlined. However, a significant challenge lies in the complexity of managing and maintaining the distributed nodes and orchestrating the computation across them, requiring considerable effort from the user. The literature presents several methodologies to address this distributed deployment paradigm. Some strategies focus on dynamic resource allocation and intelligent task distribution such as SpeziLLM (zagar2025dynamic) and AI Flow (shao2025ai). SpeziLLM is an open-source framework that dynamically distributes LLM inference across decentralised fog and edge layers, in healthcare applications. By abstracting orchestration tasks like node selection, model placement, and task splitting, SpeziLLM simplifies the integration of LLMs into mobile and healthcare environments. It prioritizes the execution of sensitive data tasks on trusted local or fog nodes while offloading less critical or computationally intensive tasks to the cloud when necessary, balancing privacy, cost, and user experience through seamless model migration, fault tolerance, and flexible scaling. AI Flow redefines communication by focusing on ”intelligence flow” rather than raw data transfer, adapting to dynamic network conditions to optimise inference across devices, edge nodes, and cloud servers. Instead of transmitting raw data, AI Flow sends only critical extracted features, significantly reducing communication overhead. It adaptively assigns portions of the inference task based on available computational resources, network bandwidth, and real-time conditions, ensuring low-latency responses and efficient model execution even in fluctuating environments. Other strategies work on collaborative inference among edge devices such as Distributed Mixture-of-Agents (MoA) (mitra2024distributed). This architecture enables multiple edge devices, each hosting a localised LLM, to collaborate through decentralised gossip protocols, achieving high-quality responses without a centralised server. Each device can independently process prompts and share intermediate results with neighboring devices using decentralized gossip algorithms. Devices act as ”proposers” generating answers and ”aggregators” refining or selecting the best response. This distributed setup enhances robustness, reduces latency, and improves answer quality compared to relying on a single device, while also ensuring queue stability despite resource limitations and varying workloads. Another approach is using the edge devices to model partitioning and distribution, cooperating as a single high performance edge device. EdgeShard (zhang2024edgeshard) is a method that facilitates efficient LLM inference by partitioning large models into smaller ”shards” and distributing them across multiple edge devices. It intelligently selects devices and allocates model parts based on their computing power, memory, and network conditions, leveraging collaborative edge computing to reduce latency, bandwidth usage, and privacy risks. By employing dynamic programming algorithms for optimized device selection and task scheduling, EdgeShard enables large models like Llama2-70B to run efficiently even in heterogeneous, resource-limited environments, significantly improving inference speed and throughput without compromising model accuracy. In (zhao2024edge) the edges are used for a cooperative inference with terminal devices. In this framework efficient LLM inference is enabled by promoting collaboration between the user’s device (terminal) and a nearby edge server. The terminal device rapidly generates speculative tokens using a lightweight model, while the edge server concurrently verifies and corrects them using a larger, more accurate LLM. This serial-parallel approach significantly reduces token generation delay and energy consumption by balancing the computational load between local and edge resources without heavy reliance on the cloud. An optimization algorithm manages model approximation and token generation to minimise delay and maintain high accuracy. Finally in distributed scenarios, guaranteeing service despite intermittent connectivity is crucial. In (zimmer2024mixture) the authors address this problem with a novel ”Mixture of Attentions” architecture for speculative decoding. This method significantly enhances a local small model’s autonomous prediction by integrating Layer Self-Attention and Cross-Attention, effectively using LLM activations when available, yet maintaining accuracy when disconnected. This approach boosts robustness, enabling continuous, accurate inference without persistent network access.